# Solving the Poisson equation using coupled Markov chains

### v3

@@ 3.3 Subsampled unbiased estimators
To prepare for the description of the proposed asymptotic variance estimator, we consider again the signed measure $\hat{\pi}$ in (3.5) and introduce measures obtained by subsampling from it. This will be used to obtain estimators with lower computational cost, for example if evaluations of $h$ are expensive in comparison to the simulation of $\hat{\pi}$. The computational benefits of subsampling in this context are related to the thinning ideas in Owen (2017).
The following result is established in Appendix B.4 as a simplification of Proposition 45. It demonstrates that the sufficient conditions for lack-of-bias and finite $p$th moments are identical for both $H_{k:\ell}$ and the subsampled estimator $S_{R}$ for any $R\geq 1$.
@@
Under Assumption 2, let $h\in L^{m}(\pi)$ for some $m>\kappa/(\kappa-1)$, ${\rm d}\pi_{0}/{\rm d}\pi\leq M$, $k,\ell\in\mathbb{N}$ with $k\leq\ell$, and $\hat{\pi}=\sum_{i=1}^{N}\omega_{i}\delta_{Z_{i}}$ be the unbiased signed measure associated with $H_{k:\ell}$, defined in (3.5). Define for some integer $R\geq 1$,
@@
where $I_{1},\ldots,I_{R}$ are conditionally independent ${\rm Categorical}\{\xi_{1},\ldots,\xi_{N}\}$ random variables with
@@
for some constants $0<a\leq b<\infty$ independent of $N$. Then $\mathbb{E}[S_{R}]=\pi(h)$ and for $p\geq 1$ such that $\frac{1}{p}>\frac{1}{m}+\frac{1}{\kappa}$, $\mathbb{E}\left[\left|S_{R}\right|^{p}\right]<\infty$.
@@
The more detailed Proposition 45 in Appendix B.4 provides a bound on $\mathbb{E}\left[\left|S_{R}\right|^{p}\right]$ but this bound does not depend on the value of $R$. For $p=2$, we can see that increasing $R$ decreases the variance of $S_{R}$, through the law of total variance
${\rm var}\left(S_{R}\right)$ $=\mathbb{E}\left[{\rm var}(S_{R}\mid\hat{\pi})\right]+{\rm var}\left(\mathbb{E}\left[S_{R}\mid\hat{\pi}\right]\right)$
$=\frac{1}{R}\mathbb{E}\left[{\rm var}(S_{1}\mid\hat{\pi})\right]+{\rm var}(\hat{\pi}(h)).$
@@ 4.1 Standard methods and motivation
The asymptotic variance $v(P,h)$ in (1.1) is commonly estimated using batch means, spectral variance (see, e.g., Flegal & Jones, 2010; Vats et al., 2018, 2019; Chakraborty et al., 2022), or initial sequence estimators (Geyer, 1992; Berg & Song, 2023). We propose new estimators of $v(P,h)$, one of which has the distinctive property of being unbiased, computable in a (random) finite time, and with $p$ finite moments under conditions given in Theorem 23. Importantly, if their two first moments are finite, unbiased estimators of $v(P,h)$ can be generated independently in parallel, and their average converges to $v(P,h)$ at the Monte Carlo rate. This compares favorably to the rate of standard estimators of $v(P,h)$.
Indeed, with batch means and spectral variance estimators, the length of the chain $t$ is decomposed as a product $a_{t}b_{t}$. For batch means, $a_{t}$ refers to a number of batches and $b_{t}$ to a batch size. For spectral variance estimators, $b_{t}$ refers to a truncation or a bandwidth parameter and $a_{t}$ denotes $t/b_{t}$. In either case, $b_{t}$ must grow with $t$ for the bias to disappear as $t\to\infty$, thus $a_{t}$ cannot grow linearly in $t$. On the other hand, the variance of these estimators is typically of the order of $a_{t}^{-1}$. Thus, even with optimal choices for $b_{t}$, the mean squared error vanishes slower than the Monte Carlo rate $t^{-1}$. Section D provides a review of relevant results from the literature on asymptotic variance estimation. This difference in convergence rates, which we observe in numerical experiments in Section 5.1, does not mean that the proposed estimators are always preferable. For some fixed computational budget, the variance of the proposed estimators can be much larger than that of classical estimators, as we observe in Section 5.2.
Asymptotic variance estimators are routinely employed to construct confidence intervals in MCMC settings, using the CLT (1.1); although Atchadé (2016) shows that consistent estimators of $v(P,h)$ may not be necessary for this task. The estimators proposed below require couplings of Markov chains, that can be used to obtain unbiased MCMC estimators as in Section 3.2. Confidence intervals for unbiased MCMC can be constructed without estimating $v(P,h)$. Nevertheless, estimation of $v(P,h)$ remains a critical task. It enables for example efficiency comparisons between unbiased and ergodic average MCMC estimators, as argued in Section 3.2. It may also be useful in comparing MCMC algorithms, as we illustrate in the numerical experiments of Section 5.2, or in choosing the tuning parameters $k$, $L$ and $\ell$, and in designing the coupled transition kernels.
@@ 4.2 Ergodic Poisson asymptotic variance estimator
where $v(\pi,h)$ is the variance of $h(X)$ under $X\sim\pi$. Consider an MCMC approach to estimate (4.1). Expectations with respect to $\pi$ can be consistently estimated from long MCMC runs. Using $t$ steps after burn-in we define the empirical measure $\pi^{\text{MC}}=t^{-1}\sum_{s=0}^{t-1}\delta_{X_{s}}$. We approximate $\pi(h)$ and $v(\pi,h)$ using the empirical mean and variance, denoted by $\pi^{\text{MC}}(h)$ and $v^{\text{MC}}(h)$.
@@
We can also modulate the relative cost of estimating the fishy function evaluations in (4.2) by generating $G(X_{s})$ only at times $s$ such that $s\mod D=0$ for some $D\in\mathbb{N}$. Preliminary runs can provide an estimate of the average cost $C$ of generating $G$ at states that approximately follow $\pi$, in units of transitions from $P$. Setting $D$ as the nearest integer to $C$ ensures that the effort of computing (4.2) is split approximately equally between the MCMC run and fishy function estimators.
In Appendix B.7, we show that the estimator $\hat{v}_{E}(P,h)$ is strongly consistent as $t\to\infty$, and satisfies a $\sqrt{t}$-CLT under Assumption 2 and moment assumptions on $h$. The following summarizes Proposition 55 and Theorem 56. An expression for the asymptotic variance can be extracted from the proof, which depends implicitly on the coupling used to define $G$.
@@
    The estimator (4.2) with $G=G_{y}$, satisfies $\hat{v}(P,h)\to_{{\rm a.s.}}v(P,h)$ as $t\to\infty$.
@@
The last item of Theorem 22 is remarkable because it implies, at least under some conditions, a rate of convergence faster than that of batch means and spectral variance estimators (Appendix D). On the other hand, EPAVE requires the implementation of successful couplings, as well as the length $t$ going to infinity. We next propose an unbiased estimator of $v(P,h)$, so that, at least, long runs can be avoided: independent copies can be generated in parallel, and averaged. Note that EPAVE is biased only because the MCMC chain does not start at stationarity. Batch means and spectral variance estimators computed in finite time would be biased even if the chains started at stationarity.
@@ 4.3 Unbiased Poisson asymptotic variance estimator
Starting again from (4.1), we propose an unbiased estimator of $v(P,h)$ by combining unbiased estimators $G(x)$ of $g(x)$ with unbiased approximations of $\pi$, as retrieved in Section 3. We thus assume that we can generate random signed measures as in (3.5), here denoted by $\hat{\pi}=\sum_{n=1}^{N}\omega_{n}\delta_{Z_{n}}$, where $N$ is a random integer, $(\omega_{n})$ are $\mathbb{R}$-valued random weights, and $(Z_{n})$ are $\mathbb{X}$-valued random variables. Recall that $\hat{\pi}$ is such that $\mathbb{E}[\hat{\pi}(h)]=\pi(h)$ for a class of test functions $h$. Combining these measures $\hat{\pi}$ with $G(x)$ in (2.4), each term in (4.1) can be estimated without bias, as we next describe. First, unbiased estimators of $v(\pi,h)=\mathbb{V}_{\pi}[h(X)]$ can be obtained using two independent unbiased measures $\hat{\pi}^{(1)}$ and $\hat{\pi}^{(2)}$, by computing
@@
Unbiased estimation of the term $\pi((h-\pi(h))\cdot g)$ in (4.1) is more involved. We first provide an informal reasoning that motivates the proposed estimator given below in (4.4) and described in pseudo-code in Algorithm 4.1. Consider the problem of estimating $\pi(h\cdot g)$ without bias, and assume that we can generate unbiased measures $\hat{\pi}=\sum_{n=1}^{N}\omega_{n}\delta_{Z_{n}}$ of $\pi$ and estimators $G(x)$ with expectation equal to $g(x)$ for all $x$. Then we can generate $\sum_{n=1}^{N}\omega_{n}h(Z_{n})\cdot G(Z_{n})$, where all $(G(Z_{n}))_{n=1}^{N}$ are conditionally independent of one another given $(Z_{n})$. Conditioning on $\hat{\pi}$, we have
@@
and then taking the expectation with respect to $\hat{\pi}$ yields $\pi(h\cdot g)$, under adequate assumptions on $h\cdot g$. However, the variable $\sum_{n=1}^{N}\omega_{n}h(Z_{n})\cdot G(Z_{n})$ requires estimators of the fishy function for all $N$ locations, and $N$ could be large. Alternatively, after generating $\hat{\pi}$ we can sample an index $I\in\{1,\ldots,N\}$ according to a Categorical distribution with strictly positive probabilities $\xi=(\xi_{1},\ldots,\xi_{N})$, and given $Z_{I}$ we can generate $G(Z_{I})$. This amounts to subsampling as described in Section 3.3. By default, we set $\xi_{k}=1/N$ for all $k\in\{1,\ldots,N\}$. Then we observe that, conditioning on $\hat{\pi}$, integrating out the randomness in $G(Z_{I})$ given $Z_{I}$, and then the randomness in $I$,
$\mathbb{E}\left[\left.\omega_{I}\frac{h(Z_{I})G(Z_{I})}{\xi_{I}}\right|\hat{\pi}\right]$ $=\mathbb{E}\left[\left.\mathbb{E}\left[\left.\omega_{I}\frac{h(Z_{I})G(Z_{I})}{\xi_{I}}\right|I,\hat{\pi}\right]\right|\hat{\pi}\right]$
$=\mathbb{E}\left[\left.\omega_{I}\frac{h(Z_{I})g(Z_{I})}{\xi_{I}}\right|\hat{\pi}\right]$
$=\sum_{k=1}^{N}\omega_{k}\frac{h(Z_{k})g(Z_{k})}{\xi_{k}}\xi_{k}$
$=\hat{\pi}(h\cdot g),$
and therefore $\omega_{I}h(Z_{I})\cdot G(Z_{I})/\xi_{I}$ is an unbiased estimator of $\pi(h\cdot g)$ that requires only one estimation of $g$ at $Z_{I}$. The estimator proposed below employs $R\geq 1$ estimators of the fishy function for each signed measure $\hat{\pi}$, where $R$ is a tuning parameter. Its choice and the selection probabilities $\xi$ are discussed in Section 4.4.
@@
      Draw $I^{(r,j)}$ among $\{1,\ldots,N^{(j)}\}$ with probabilities $(\xi_{1}^{(j)},\ldots,\xi_{N^{(j)}}^{(j)})$, for $r\in\{1,\ldots,R\}$.
@@
      Evaluate $h(Z_{I^{(r,j)}}^{(j)})$ and generate estimator $G(Z_{I^{(r,j)}}^{(j)})$ in (2.4), for $r\in\{1,\ldots,R\}$.
@@
We gather the above considerations to define the proposed estimator. We write $\hat{\pi}^{(j)}$ as $\sum_{n=1}^{N^{(j)}}\omega_{n}^{(j)}\delta_{Z_{n}^{(j)}}$ for $j\in\{1,2\}$. Given $\hat{\pi}^{(j)}$, we sample integers $I^{(r,j)}\in\{1,\ldots,N^{(j)}\}$ with probabilities $(\xi_{1}^{(j)},\ldots,\xi_{N^{(j)}}^{(j)})$, independently for $r\in\{1,\ldots,R\}$. Noting that each $(h(x)-\pi(h))g(x)$ is the expectation of $(h(x)-\hat{\pi}^{(j)}(h))G(x)$ given $x$, we obtain
$\pi((h-\pi(h))\cdot g)=\mathbb{E}\left[\frac{1}{2R}\sum_{i\neq j\in\{1,2\}}\sum_{r=1}^{R}\frac{\omega_{I^{(r,j)}}^{(j)}}{\xi_{I^{(r,j)}}^{(j)}}(h(Z_{I^{(r,j)}}^{(j)})-\hat{\pi}^{(i)}(h))G(Z_{I^{(r,j)}}^{(j)})\right].$
@@
$\hat{v}(P,h)=-\hat{v}(\pi,h)+\frac{1}{R}\sum_{i\neq j\in\{1,2\}}\sum_{r=1}^{R}\frac{\omega_{I^{(r,j)}}^{(j)}}{\xi_{I^{(r,j)}}^{(j)}}(h(Z_{I^{(r,j)}}^{(j)})-\hat{\pi}^{(i)}(h))G(Z_{I^{(r,j)}}^{(j)}),$ (4.4)
and its generation is described in Algorithm 4.1. The total cost of $\hat{v}(P,h)$ will typically be dominated by the cost of obtaining two signed measures $\hat{\pi}^{(j)}$ for $j\in\{1,2\}$ and $2R$ estimators of evaluations of $g$. We call $\hat{v}(P,h)$ the unbiased Poisson asymptotic variance estimator (UPAVE).
We show that under Assumption 2, UPAVE is unbiased and has $p$ finite moments whenever $h$ has sufficiently many moments. The following statement combines Theorem 53 and Remark 54 in Appendix B.6, with the further assumption that $\xi_{k}^{(j)}=1/N^{(j)}$ for all $k\in\{1,\ldots,N^{(j)}\}$, for simplicity.
@@
Under Assumption 2, let $h\in L^{m}(\pi)$ for some $m>2\kappa/(\kappa-2)$. Assume $\xi_{k}^{(j)}=1/N^{(j)}$ for $k\in\{1,\ldots,N^{(j)}\}$. Then for any $R\geq 1$ and $\pi$-almost all $y$, $\mathbb{E}\left[\hat{v}(P,h)\right]=v(P,h)$ and for $p\geq 1$ such that $\frac{1}{p}>\frac{2}{m}+\frac{2}{\kappa}$, $\mathbb{E}\left[\left|\hat{v}(P,h)\right|^{p}\right]<\infty$.

### v4

@@ 3.3 Subsampled unbiased estimators
To prepare for the description of the estimators in Section 4, we simplify the notation of the signed measure $\hat{\pi}_{k:\ell}$ in (3.5)-(3.6). Let $N=(\ell-k+1)+2\max(0,\tau-(k+L))$ be the random number of atoms in $\hat{\pi}_{k:\ell}$, and denote these atoms by $Z_{1},\ldots,Z_{N}$. The first $\ell-k+1$ atoms correspond to $X_{k},\ldots,X_{\ell}$, and the following $2\max(0,\tau-(k+L))$ atoms (if any) correspond to $X_{k+L},Y_{k},\ldots,X_{\tau-1},Y_{\tau-1-L}$. There may be duplicate atoms among $Z_{1},\ldots,Z_{N}$. We can write
$\hat{\pi}_{k:\ell}({\rm d}x)=\sum_{i=1}^{N}\omega_{i}\delta_{Z_{i}}({\rm d}x),$ (3.7)
where $\omega_{i}$ are the weights, of the form $(\ell-k+1)^{-1}$ for the first $\ell-k+1$ atoms, and of the form ±$\pm v_{t}(\ell-k+1)^{-1}$ with $v_{t}$ defined in (3.6) and $t\in\{k+L,\ldots,\tau-1\}$ for the remaining atoms. We further simplify the notation by writing $\hat{\pi}$ instead of $\hat{\pi}_{k:\ell}$. For a function $h$, $\hat{\pi}(h)=\sum_{i=1}^{N}\omega_{i}h(Z_{i})=H_{k:\ell}$, where $H_{k:\ell}$ is the unbiased MCMC estimator in Definition 16.
We can obtain estimators of $\pi(h)$ by subsampling the atoms of $\hat{\pi}$ in (3.7). This results in estimators with lower computational cost, for example if evaluations of the test function $h$ are expensive in comparison to the simulation of $\hat{\pi}$. The computational benefits of subsampling in this context are related to the thinning ideas in Owen (2017). The following result is established in Appendix B.4. It demonstrates that the sufficient conditions for lack-of-bias and finite $p$th moments are identical for $H_{k:\ell}$ and the subsampled estimator $S_{R}$ for any $R\geq 1$.
@@
Under Assumption 2, let $h\in L^{m}(\pi)$ for some $m>\kappa/(\kappa-1)$, ${\rm d}\pi_{0}/{\rm d}\pi\leq M$, $k,\ell\in\mathbb{N}$ with $k\leq\ell$, and $\hat{\pi}=\sum_{i=1}^{N}\omega_{i}\delta_{Z_{i}}$ be the unbiased signed measure in (3.7). Define for some integer $R\geq 1$,
@@
where $I_{1},\ldots,I_{R}$ are conditionally independent ${\rm Categorical}\{\xi_{1},\ldots,\xi_{N}\}$ variables with
@@
for some constants $0<a\leq b<\infty$ that may be functions of $k,\ell,L$ but not $\tau$. Then $\mathbb{E}[S_{R}]=\pi(h)$ and for $p\geq 1$ such that $\frac{1}{p}>\frac{1}{m}+\frac{1}{\kappa}$, $\mathbb{E}\left[\left|S_{R}\right|^{p}\right]<\infty$.
@@
The more detailed Proposition B.22 in Appendix B.4 provides a bound on $\mathbb{E}\left[\left|S_{R}\right|^{p}\right]$ but this bound does not depend on the value of $R$. For $p=2$, we can see that increasing $R$ decreases the variance of $S_{R}$, through the law of total variance
${\rm var}\left(S_{R}\right)$ $=\mathbb{E}\left[{\rm var}(S_{R}\mid\hat{\pi})\right]+{\rm var}\left(\mathbb{E}\left[S_{R}\mid\hat{\pi}\right]\right)=\frac{1}{R}\mathbb{E}\left[{\rm var}(S_{1}\mid\hat{\pi})\right]+{\rm var}(\hat{\pi}(h)).$
@@ 4.1 Standard methods and motivation
The asymptotic variance $v(P,h)$ in (1.1) is commonly estimated using batch means, spectral variance (see, e.g., Flegal & Jones, 2010; Vats et al., 2018, 2019; Chakraborty et al., 2022), or initial sequence estimators (Geyer, 1992; Berg & Song, 2023). We propose new estimators of $v(P,h)$, one of which has the distinctive property of being unbiased, computable in a (random) finite time, and with $p$ finite moments under conditions given in Theorem 23. Importantly, if their two first moments are finite, unbiased estimators of $v(P,h)$ can be generated independently in parallel, and their average converges at the Monte Carlo rate. This compares favorably to the rate of standard estimators of $v(P,h)$.
Indeed, with batch means and spectral variance estimators, the length of the chain $t$ is decomposed as a product $a_{t}b_{t}$. For batch means, $a_{t}$ refers to a number of batches and $b_{t}$ to a batch size. For spectral variance estimators, $b_{t}$ refers to a truncation or bandwidth parameter and $a_{t}$ denotes $t/b_{t}$. In either case, $b_{t}$ must grow with $t$ for the bias to vanish as $t\to\infty$, thus $a_{t}$ cannot grow linearly in $t$. On the other hand, the variance of these estimators is typically of the order of $a_{t}^{-1}$. Thus, even with optimal choices for $b_{t}$, the mean squared error vanishes slower than the Monte Carlo rate $t^{-1}$. Appendix C provides a review of relevant results on asymptotic variance estimation. This difference in convergence rates, which we observe in numerical experiments in Section 5.1, does not mean that the proposed estimators are always preferable. For some fixed computational budget, the variance of the proposed estimators can be much larger than that of classical estimators, as we observe in Section 5.2.
Asymptotic variance estimators are routinely employed to construct confidence intervals in MCMC settings, using the CLT (1.1); although Atchadé (2016) shows that consistent estimators of $v(P,h)$ may not be necessary for this task. The estimators proposed below require couplings of Markov chains, that can be used to obtain unbiased MCMC estimators as in Section 3.2. Confidence intervals for unbiased MCMC can be constructed without estimating $v(P,h)$. Nevertheless, estimation of $v(P,h)$ remains a critical task. It enables for example efficiency comparisons between unbiased and ergodic average MCMC estimators, as argued in Section 3.2. It may also be useful in comparing MCMC algorithms, as we illustrate in the numerical experiments of Section 5.2, or in choosing the tuning parameters $k$, $L$ and $\ell$.
@@ 4.2 Ergodic Poisson asymptotic variance estimator
where $v(\pi,h)$ is the variance of $h(X)$ under $X\sim\pi$. Consider an MCMC approach to estimate (4.1). Expectations with respect to $\pi$ can be consistently estimated from long MCMC runs. Using $t$ steps of a Markov chain evolving with transition $P$, after a burn-in period, and re-indexing iterations so that $t=0$ corresponds to the first iteration after the discarded burn-in period, we define the empirical measure $\pi^{\text{MC}}=t^{-1}\sum_{s=0}^{t-1}\delta_{X_{s}}$. We approximate $\pi(h)$ and $v(\pi,h)$ using the empirical mean and variance, denoted by $\pi^{\text{MC}}(h)$ and $v^{\text{MC}}(h)$.
@@
We can modulate the relative cost of estimating the fishy function evaluations in (4.2) by generating $G(X_{s})$ at times $s$ such that $s\mod D=0$ for $D\in\mathbb{N}$. Preliminary runs can provide an estimate of the average cost $C$ of generating $G$ at states that approximately follow $\pi$, in units of transitions from $P$. Setting $D$ as $\lceil C\rceil$ ensures that the effort of computing (4.2) is split approximately equally between the MCMC run and the estimators $G$.
In Appendix B.7, we show that the estimator $\hat{v}_{E}(P,h)$ is strongly consistent as $t\to\infty$, and satisfies a $\sqrt{t}$-CLT under Assumption 2 and moment assumptions on $h$. The following summarizes Proposition B.32 and Theorem B.33. An expression for the asymptotic variance of EPAVE can be extracted from the proof.
@@
    The estimator (4.2) with $G=G_{y}$, satisfies $\hat{v}_{E}(P,h)\to_{{\rm a.s.}}v(P,h)$ as $t\to\infty$.
@@
The last item of Theorem 22 is remarkable because it implies, at least under some conditions, a rate of convergence faster than that of batch means and spectral variance estimators, see Appendix C. On the other hand, EPAVE requires the implementation of successful couplings, as well as the length $t$ going to infinity. We next propose an unbiased estimator of $v(P,h)$, so that, at least, long runs can be avoided: independent copies can be generated in parallel, and averaged. Note that EPAVE is biased only because the MCMC chain does not start at stationarity. Batch means and spectral variance estimators computed in finite time would be biased even if the chains started at stationarity.
@@ 4.3 Unbiased Poisson asymptotic variance estimator
Starting again from (4.1), we propose an unbiased estimator of $v(P,h)$ by combining unbiased estimators $G(x)$ of $g(x)$ with unbiased approximations of $\pi$, as retrieved in Section 3. We thus assume that we can generate random signed measures as in (3.5), denoted by $\hat{\pi}=\sum_{n=1}^{N}\omega_{n}\delta_{Z_{n}}$ as in (3.7). Recall that $\hat{\pi}$ is such that $\mathbb{E}[\hat{\pi}(h)]=\pi(h)$ for a class of test functions $h$. Combining these measures $\hat{\pi}$ with $G(x)$ in (2.4), each term in (4.1) can be estimated without bias, as we next describe. First, unbiased estimators of $v(\pi,h)=\mathbb{V}_{\pi}[h(X)]$ can be obtained using two independent unbiased measures $\hat{\pi}^{(1)}$ and $\hat{\pi}^{(2)}$, by computing
@@
Unbiased estimation of the term $\pi((h-\pi(h))\cdot g)$ in (4.1) is more involved. We first provide an informal reasoning that motivates the proposed estimator given below in (4.4) and described in pseudocode in Algorithm 4.1. Consider the problem of estimating $\pi(h\cdot g)$ without bias, and assume that we can generate unbiased measures $\hat{\pi}$ of $\pi$ and estimators $G(x)$ with expectation equal to $g(x)$ for all $x$. Then we can generate $\sum_{n=1}^{N}\omega_{n}h(Z_{n})\cdot G(Z_{n})$, where all $(G(Z_{n}))_{n=1}^{N}$ are conditionally independent given $(Z_{n})$. Conditioning on $\hat{\pi}$, we have
@@
and then taking the expectation with respect to $\hat{\pi}$ yields $\pi(h\cdot g)$, under adequate assumptions on $h\cdot g$. However, the variable $\sum_{n=1}^{N}\omega_{n}h(Z_{n})\cdot G(Z_{n})$ requires estimators of the fishy function for all $N$ locations, and $N$ could be large. Alternatively, after generating $\hat{\pi}$ we can sample an index $I\in\{1,\ldots,N\}$ according to a Categorical distribution with strictly positive probabilities $\xi=(\xi_{1},\ldots,\xi_{N})$, and given $Z_{I}$ we can generate $G(Z_{I})$. This amounts to subsampling as described in Section 3.3. By default, we set $\xi_{n}=1/N$ for all $n\in\{1,\ldots,N\}$. Then we observe that, conditioning on $\hat{\pi}$, integrating out the randomness in $G(Z_{I})$ given $Z_{I}$, and then the randomness in $I$,
$\mathbb{E}\left[\left.\omega_{I}\xi_{I}^{-1}h(Z_{I})G(Z_{I})\right|\hat{\pi}\right]$ $=\mathbb{E}\left[\left.\mathbb{E}\left[\left.\omega_{I}\xi_{I}^{-1}h(Z_{I})G(Z_{I})\right|I,\hat{\pi}\right]\right|\hat{\pi}\right]$
$=\mathbb{E}\left[\left.\omega_{I}\xi_{I}^{-1}h(Z_{I})g(Z_{I})\right|\hat{\pi}\right]=\sum_{n=1}^{N}\omega_{n}h(Z_{n})g(Z_{n})=\hat{\pi}(h\cdot g),$
and therefore $\omega_{I}\xi_{I}^{-1}h(Z_{I})\cdot G(Z_{I})$ is an unbiased estimator of $\pi(h\cdot g)$ that requires only one estimation of $g$ at $Z_{I}$. The estimator proposed below employs $R\geq 1$ estimators of the fishy function for each signed measure $\hat{\pi}$, where $R$ is a tuning parameter. Its choice and the selection probabilities $\xi$ are discussed in Section 4.4.
@@
      Draw $I^{(j)}_{r}$ among $\{1,\ldots,N^{(j)}\}$ with probabilities $(\xi_{1}^{(j)},\ldots,\xi_{N^{(j)}}^{(j)})$, for $r\in\{1,\ldots,R\}$.
@@
      Evaluate $h(Z_{I^{(j)}_{r}}^{(j)})$ and generate estimator $G(Z_{I^{(j)}_{r}}^{(j)})$ in (2.4), for $r\in\{1,\ldots,R\}$.
@@
We gather the above considerations to define the proposed estimator. We write $\hat{\pi}^{(j)}$ as $\sum_{n=1}^{N^{(j)}}\omega_{n}^{(j)}\delta_{Z_{n}^{(j)}}$ for $j\in\{1,2\}$. Given $\hat{\pi}^{(j)}$, we sample integers $I^{(j)}_{r}\in\{1,\ldots,N^{(j)}\}$ with probabilities $(\xi_{1}^{(j)},\ldots,\xi_{N^{(j)}}^{(j)})$, independently for $r\in\{1,\ldots,R\}$. Noting that each $(h(x)-\pi(h))g(x)$ is the expectation of $(h(x)-\hat{\pi}^{(j)}(h))G(x)$ given $x$, we obtain
$\pi((h-\pi(h))\cdot g)=\mathbb{E}\left[\frac{1}{2R}\sum_{i\neq j\in\{1,2\}}\sum_{r=1}^{R}\frac{\omega_{I^{(j)}_{r}}^{(j)}}{\xi_{I^{(j)}_{r}}^{(j)}}(h(Z_{I^{(j)}_{r}}^{(j)})-\hat{\pi}^{(i)}(h))G(Z_{I^{(j)}_{r}}^{(j)})\right].$
@@
$\hat{v}(P,h)=-\hat{v}(\pi,h)+\frac{1}{R}\sum_{i\neq j\in\{1,2\}}\sum_{r=1}^{R}\frac{\omega_{I^{(j)}_{r}}^{(j)}}{\xi_{I^{(j)}_{r}}^{(j)}}(h(Z_{I^{(j)}_{r}}^{(j)})-\hat{\pi}^{(i)}(h))G(Z_{I^{(j)}_{r}}^{(j)}),$ (4.4)
and its generation is described in Algorithm 4.1. The cost of $\hat{v}(P,h)$ will typically be dominated by the cost of obtaining $\hat{\pi}^{(j)}$ for $j\in\{1,2\}$ and $2R$ estimators of evaluations of $g$. We call $\hat{v}(P,h)$ the unbiased Poisson asymptotic variance estimator (UPAVE).
We show that under Assumption 2, UPAVE is unbiased and has $p$ finite moments whenever $h$ has sufficiently many moments. The following statement combines Theorem B.30 and Remark B.31 in Appendix B.6, with the further assumption that $\xi_{n}^{(j)}=1/N^{(j)}$ for all $n\in\{1,\ldots,N^{(j)}\}$, for simplicity.
@@
Under Assumption 2, let $h\in L^{m}(\pi)$ for some $m>2\kappa/(\kappa-2)$, and ${\rm d}\pi_{0}/{\rm d}\pi\leq M$. Assume $\xi_{n}^{(j)}=1/N^{(j)}$ for $n\in\{1,\ldots,N^{(j)}\}$. Then for any $R\geq 1$ and $\pi$-almost all $y$, $\mathbb{E}\left[\hat{v}(P,h)\right]=v(P,h)$ and for $p\geq 1$ such that $\frac{1}{p}>\frac{2}{m}+\frac{2}{\kappa}$, $\mathbb{E}\left[\left|\hat{v}(P,h)\right|^{p}\right]<\infty$.